a boat takes time t to go downstream from point A to point B  and time 2t to go upstream from point B to point A.If the speed of the stream is v and the speed of the boat with respect to the stream is u,then which of the following gives correct relationship b/w u & v?


a)u=v


b)u=3v


c)v=2u


d)v=3u


choose and explain?

Grade:12

1 Answers

Ashwin Muralidharan IIT Madras
10 years ago

Hi Pratiksha,

 

While going DS, the speed of the boat will be (u+v) wrt ground

and going US it would be (u-v).

Let x be the distance between the points A and B.

 

So x/(u+v) = t

and x/(u-v) = 2t

 

or 2x/(u+v) = x/(u-v).

or u=3v.

Option (B)

 

Regards,

Ashwin (IIT Madras).